OpenStudy (anonymous):

Hi can someone please help me solve one algebra question? Solve: 2x + 1 >= 5 or -3x > -9 Thank you in advance(:

OpenStudy (wolfe8):

Treat the inequality as you would a normal equation. Can you do it from there?

OpenStudy (nurali):

2x + 1 >= 5 2x>= 5-1 2x>=4 x>=4/2 x>=2

OpenStudy (nurali):

-3x > -9 -x>-9/3 -x>-3 x<3

OpenStudy (anonymous):

Nurali thank you so much! can you help me with this one to? its the last one. -4y+17<=29

OpenStudy (nurali):

My Pleasure.

OpenStudy (nurali):

-4y+17<=29 -4y<=29-17 -4y<=12 y<=-12/4 y<=-3

OpenStudy (anonymous):

(: Is there a way that I can give you more medals?